2) If someone tells you that they prefer watching the 49ers to watching the Sharks, ro prefer watching the Sharks to watching the Warriors, and prefer watching the Warriors to watching the 49ers, what basic assumption of economics are they violating? 3) Suppose demand in the market for oranges is characterized by the demand equation QD 25-0.25P. a) What is the marginal revenue of the 5th orange produced? b) What is the marginal revenue of the 10th orange produced?
Answer #1

Answer-2 The assumption of transitivity is getting voilated.

Transitivity assumption says that when there are bundles: x,y and z.

If x is preferred over y, y is preferred over x then x will be preferred over z.

But here in this case, watching warriors is preferred over 49ers. To not voilate transitivity, watching 49ers must be preferred over watching warriors.

Answer-3 Q= 25-0.25P

P = (25-Q)/0.25

P = 100- 4 Q

Total Revenue = P.Q

= (100-4Q)Q

= 100Q -4Q^2

Marginal revenue = dTR/dQ = 100-8Q

a) Marginal revenue of 5th orange is 100-8(5) = 60

b) Marginal revenue of 10th orange is 100-8(10) = 20